# DESY-25-144

### v1

@@ Abstract
Cosmological phase transitions are a frequent phenomenon in particle physics models beyond the Standard Model, and the corresponding gravitational wave signal offers a key probe of new physics in the early Universe. Depending on the underlying microphysics, the transition can exhibit either direct or inverse hydrodynamics, leading to a different phenomenology. Most studies to date have focused on direct transitions, where the cosmic fluid is pushed or dragged by the expanding vacuum bubbles. In contrast, inverse phase transitions are characterized by fluid profiles where the plasma is sucked in by the expanding bubbles. Using the sound shell model, we derive and compare the gravitational wave spectra from sound waves for direct and inverse phase transitions, providing new insights into the potential observable features and the possibility of discriminating among the various fluid solutions in gravitational wave experiments.
@@ I Introduction
We study cosmological phase transitions (PTs) in the early Universe, with a focus on first-order phase transitions (FOPTs) that proceed via the nucleation and expansion of bubbles of the true vacuum within a sea of false vacuum; these transitions can trigger baryogenesis [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16], lead to the production of dark matter [17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28], generate primordial black holes [29, 30, 31, 32, 33], and source stochastic gravitational wave (GW) backgrounds [34, 35, 36, 37, 38]. The discovery prospects with space- and ground-based next-generation GW interferometers such as LISA, ET, CE, BBO, and DECIGO open a unique window onto otherwise inaccessible sectors of particle physics, as many well-motivated Beyond the Standard Model (BSM) frameworks naturally realize cosmological PTs, including composite Higgs models [39, 40, 41, 42, 43], extended Higgs sectors [44, 45, 46, 47, 48, 49, 50, 51, 52, 53], axion models [54, 55], dark sectors [56, 57, 58, 59, 60], and BLB\!-\!L-breaking sectors [61, 62].
Against this backdrop, interest in inverse phase transitions has grown, particularly in cosmological settings that feature transient heating phases — for example, reheating after inflation [63], regions influenced by the rarefaction tails of bubbles from a preceding (direct) phase transition [64], and localized heating near primordial black holes (PBHs) [65] and in the core of dense neutron stars [66, 67, 68] — in which the plasma’s thermodynamic trajectory can favor an inverse transition and thereby alter the standard, monotonic-cooling picture. The hydrodynamics of inverse phase transitions has been investigated in [69, 67], highlighting how heating backgrounds and nontrivial flow profiles modify the interplay between bubble expansion, shock/rarefaction fronts, and energy-momentum transport. Perhaps even more surprisingly, symmetry-breaking phase transitions while standard cosmological cooling have recently been shown to arise within supersymmetric models [70] and BSM realisations of the EWPT [71] [^1] Notice that the symmetry presented in [71] was strictly speaking a symmetry restoring direct phase transitions., adding theoretical impetus to revisit and systematize inverse-transition dynamics.
@@
The LTE approximation [74, 75, 76, 77] is a simple and flexible assumption customarily used to obtain an upper bound on the wall velocity. Within this approximation, the plasma on both sides and inside the bubble wall is assumed to remain in equilibrium, with entropy conservation linking the temperatures as $T_{-}/T_{+}=\gamma_{+}/\gamma_{-}$, where subscripts $\pm$’ denote quantities on the outer (‘++’) and inner (‘-’) side of the wall, and $\gamma_{\pm}$ are the boost factor of the fluid velocity on either side of the wall (in the wall frame). Under this assumption, the hydrodynamic equations reduce to a one-parameter family of solutions that can be fully determined. The matching conditions across the wall relate the fluid velocities and temperatures without having to solve the Boltzmann equation explicitly. This approximation then provides a controlled and efficient way to estimate the wall dynamics and the resulting hydrodynamic profiles driving the GW source.
The SSM provides a physically transparent description of how overlapping spherical sound shells generate a stochastic background of GWs within the approximation of linear and Gaussian dynamics. Originally proposed in Refs. [78, 79], it has been extended to include e.g. the expansion history of the Universe and different equations of state [80, 81, 82, 83] and used in observational studies [84, 85, 86, 87, 88]. In this work we adopt the SSM, closely following the implementation in Ref. [81], as our baseline framework to capture the dominant acoustic contribution for both direct and inverse transitions, keeping in mind however the qualitative features that are unique to inverse transitions, namely the inflow fluid dynamics across the bubble wall, which motivate dedicated numerical simulations as the next step for precise GW predictions.
@@ II Reminder on inverse phase transitions
Before turning to hydrodynamics, we clarify the distinction between direct and inverse first-order phase transitions in terms of the sign of the latent heat. Let $L$ denote the latent heat evaluated at the transition (with $\Delta s$ the entropy jump across the phases). Then:
  • Direct PT ($L>0$). The transition releases energy into the environment. The injected energy drives an outward pressure gradient in the plasma, sourcing fluid outflows from the advancing bubble wall, going “away” as seen from the center of the bubble.
  • Inverse PT ($L<0$). The transition absorbs energy from the environment. Energy is drawn from the surrounding plasma into the bubble interface, which tends to establish inward pressure gradients and correspondingly favors fluid motion toward the center of the bubble.
In this description, “direct” versus “inverse” refers solely to the sign of $L$: positive $L$ implies energy release; negative $L$ implies energy uptake. Because inverse transitions require extracting heat from the medium ($L<0$), they are intrinsically thermal processes and cannot occur at $T=0$.
This reversed fluid flow has profound implications for the dynamics of the transition. The hydrodynamic solutions that describe the plasma motion, such as detonations, deflagrations, and hybrids, are qualitatively modified, as the pressure and velocity profiles now correspond to an inward rather than outward propagation. Moreover, the interplay between the fluid and the scalar field becomes even more critical: while in direct transitions the potential gradient primarily drives the expansion, in inverse transitions it competes with the inward motion of the fluid, and this coupling determines the steady-state wall velocity as well as the energy budget going into bulk–fluid kinetic energy.
@@ II.1 Hydrodynamics of inverse phase transitions
beyond which either the detonation or the deflagration branch is selected and the composite construction ceases to satisfy the matching and causality constraints. By definition, [^4] First introduced in [69], $v_{J,\rm inv}$ marks the transition between inverse detonation and inverse hybrid regimes and can be viewed as the wall speed of the fastest inverse detonation (or, equivalently, the slowest inverse hybrid). $v_{J,\rm inv}$ is the boundary value at which the rarefaction and compression attachments become sonic and the hybrid solution bifurcates into a single-wave configuration.
@@ III The velocity in the LTE regime
In principle, the velocity of the bubble is not determined by the hydrodynamical properties of the bubble, and microphysics considerations need to be invoked. We can verify it by counting the number of parameters describing the phase transition, four: $T_{+},T_{-},v_{+},v_{-}$, and the number of conserving equations written down in Eq.˜3, two: conservation of the stress energy tensor. So even for a given nucleation temperature (determined by the nucleation time of the bubble), this leaves one parameter unconstrained, which we typically take to be the velocity of the wall, $\xi_{w}$. In principle, this would require studying the microphysics of the plasma around the bubble wall. In the following, we will follow an orthogonal approach.
@@ III.1 LTE framework
With this equation at hand, we can study the pressure for the direct and the inverse phase transitions. First of all, we can study this relation in the case of a just–nucleated bubble. The bubble is taken to be nucleated at rest with a vanishing wall velocity, $\xi_{w}\to 0$, and correspondingly $v_{+}=v_{-}\to 0$, consequently we have that
@@ III.2 LTE pressure for direct phase transitions
Figure 4: Phase diagram of the LTE wall velocity in the $(\alpha_{N},\Psi)$ plane. Coloured squares show the numerical wall speed. The solid black boundary is the full $\mathcal{P}_{\mathrm{LTE}}=0$ contour: its straight left segment is the static–wall limit $\mathcal{P}_{\mathrm{LTE}}(\xi_{w}=0)=0$ (a newly nucleated bubble does not expand). In the direct region, the curved branch marks the onset of runaway, equivalently $\mathcal{P}_{\mathrm{LTE}}(\xi_{w}=\xi_{J})<0$; in the inverse region, it is set by the slowest hybrid just past the inverse Jouguet speed, $\mathcal{P}_{\mathrm{LTE}}(\xi_{w}=\xi_{J,\mathrm{inv}}^{+})<0$. The green band highlights parameters satisfying the runaway criterion. The dotted vertical line at $\alpha_{N}\simeq-0.082$ indicates the point at which the kinematically forbidden region opens up, and the slowest inverse hybrid solution is always a stable solution.
@@ III.3 The LTE pressure for inverse phase transitions
As such, this boundary is expressed in terms of $\alpha_{+}$. To express it in terms of $\alpha_{N}$, we then rescale
$\alpha_{+}\to\alpha_{N}\,\frac{\alpha_{+,\rm slowest}}{\alpha_{N,\rm slowest}}\,$,
and substitute into Eq.˜30 to obtain the black curve delineating the transition between LTE‐velocity solutions and true runaway.
@@ IV Sound-shell model in an expanding Universe
In this section, we compute the gravitational-wave signal from first-order phase transitions with the explicit goal of disentangling direct from inverse dynamics. Our strategy is to model the acoustic stage within the sound-shell framework, in which the velocity field produced by the expanding bubbles is described as a superposition of spherically symmetric shells whose overlap sources tensor perturbations on a Friedmann background. The approach is well-established for phase transitions of weak to intermediate strengths, from its original formulation and first quantitative predictions [78, 79] to observational applications [84, 85, 86, 87, 88], and it has been extended to include the impact of the cosmic expansion history on the build-up and time coherence of the source [80] as well as different equations of state [82]. Here we adopt the formulation that clarifies the infrared momentum dependence and the emergence of intermediate spectral slopes [81], which we implement to track how the hydrodynamic differences between direct and inverse transitions propagate into the shape and normalization of the spectrum.
@@ V.1 Spectral angles: from pure shape to amplitude-aware comparisons
We highlight a distinctive low–velocity effect: the modified spectral angle tends to flatten as the wall speed decreases. Tracing this back to the energy budget, we find that the multi–bubble kinetic fraction $\Omega_{K}$ also flattens in the $\xi_{w}\!\to\!0$ limit. A steeper suppression might have been expected by considering the single–bubble efficiency only, which is known to drop rapidly at small $\xi_{w}$ [69]. As we argue in Appendix˜B, this is due to the fact that the overlap of many bubbles at collision compensates for the poor conversion of the isolated expanding bubbles.
@@ Acknowledgements.
It is a pleasure to thank Mark Hindmarsh, Simona Procacci and Alberto Roper Pol for helpful discussions, as well as Wenyuan Ai and Benoit Laurent for fruitful comments on the manuscript. EM and GB acknowledge support from the Spanish Research Agency (Agencia Estatal de Investigación, MCIN/AEI/10.13039/501100011033) via the IFT Severo Ochoa Center of Excellence grant CEX2020-001007-S. In addition, EM is supported through the grants CNS2023-14453 and PID2022-137127NB-I00 funded by MCIN/AEI/10.13039/501100011033, the European Union NextGenerationEU/PRTR and ESF+, and GB is supported by the grant CNS2023-145069 funded by MICIU/AEI/10.13039/501100011033 and by the European Union NextGenerationEU/PRTR. MV is supported by the “Excellence of Science - EOS” - be.h project n.30820817, and by the Strategic Research Program High-Energy Physics of the Vrije Universiteit Brussel. SB is supported by the Deutsche Forschungsgemeinschaft under Germany’s Excellence Strategy - EXC 2121 Quantum universe - 390833306.

### v2

@@ Abstract
Cosmological first order phase transitions are a frequent phenomenon in particle physics beyond the Standard Model, and the corresponding gravitational wave signal offers a key probe of new physics in the early Universe. Depending on the underlying microphysics, the transition can exhibit either direct or inverse hydrodynamics, leading to a different phenomenology. Most studies to date have focused on direct transitions, where the cosmic fluid is pushed or dragged by the expanding vacuum bubbles. In contrast, inverse phase transitions are characterized by fluid profiles where the plasma is sucked in by the expanding bubbles. Using the sound shell model, we derive and compare the gravitational wave spectra from sound waves for direct and inverse phase transitions, providing new insights into the potential observable features and the possibility of discriminating among the various fluid solutions in gravitational wave experiments.
@@ I Introduction
We study cosmological phase transitions (PTs) in the early Universe, with a focus on first-order phase transitions (FOPTs) that proceed via the nucleation and expansion of bubbles of the true vacuum within a sea of false vacuum; these transitions can trigger baryogenesis [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16], lead to the production of dark matter [17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28], generate primordial black holes [29, 30, 31, 32, 33], and source stochastic gravitational wave (GW) backgrounds [34, 35, 36, 37, 38]. The discovery prospects with space- and ground-based next-generation GW interferometers such as LISA, ET, CE, BBO, and DECIGO open a unique window onto otherwise inaccessible sectors of particle physics, as many well-motivated Beyond the Standard Model (BSM) frameworks naturally realize cosmological PTs, including composite Higgs models [39, 40, 41, 42, 43], extended Higgs sectors [44, 45, 46, 47, 48, 49, 50, 51, 52, 53], axion models [54, 55], dark sectors [56, 57, 58, 59, 60, 61, 62], and BLB\!-\!L-breaking sectors [63, 64].
Against this backdrop, interest in inverse phase transitions has grown, particularly in cosmological settings that feature transient heating phases — for example, reheating after inflation or from energy injection [65, 66, 67, 68, 69], regions influenced by the rarefaction tails of bubbles from a preceding (direct) phase transition [70], and localized heating near primordial black holes (PBHs) [71] and in the core of dense neutron stars [72, 73, 74] — in which the plasma’s thermodynamic trajectory can favor an inverse transition and thereby alter the standard, monotonic-cooling picture. The hydrodynamics of inverse phase transitions has been investigated in [75, 73], highlighting how heating backgrounds and nontrivial flow profiles modify the interplay between bubble expansion, shock/rarefaction fronts, and energy-momentum transport. Perhaps even more surprisingly, symmetry-breaking phase transitions while standard cosmological cooling have recently been shown to arise within supersymmetric models [76] and BSM realisations of the EWPT [77], adding theoretical motivation to revisit and systematize inverse-transition dynamics.
@@
The LTE approximation [80, 81, 82, 83] is a simple and flexible assumption customarily used to obtain an estimate of the wall velocity [84]. Within this approximation, the plasma on both sides and inside the bubble wall is assumed to remain in equilibrium, with entropy conservation linking the temperatures as $T_{-}/T_{+}=\gamma_{+}/\gamma_{-}$, where subscripts $\pm$’ denote quantities on the outer (‘++’) and inner (‘-’) side of the wall, and $\gamma_{\pm}$ are the boost factor of the fluid velocity on either side of the wall (in the wall frame). Under this assumption, the hydrodynamic equations reduce to a one-parameter family of solutions that can be fully determined. The matching conditions across the wall relate the fluid velocities and temperatures without having to solve the Boltzmann equation explicitly. This approximation then provides a controlled and efficient way to estimate the wall dynamics and the resulting hydrodynamic profiles driving the GW source.
The SSM provides a physically transparent description of how overlapping spherical sound shells generate a stochastic background of GWs within the approximation of linear and Gaussian dynamics. Originally proposed in Refs. [85, 86], it has been extended to include e.g. the expansion history of the Universe and different equations of state [87, 88, 89, 90] and used in observational studies [91, 92, 93, 94, 95]. In this work we adopt the SSM, closely following the implementation in Ref. [88], as our baseline framework to capture the dominant acoustic contribution for both direct and inverse transitions. It is important to emphasize that the qualitative features unique to inverse transitions, most notably the inflow fluid dynamics across the bubble wall, call for dedicated numerical simulations as a crucial next step toward accurate gravitational-wave predictions.
@@ II Reminder on inverse phase transitions
Before turning to hydrodynamics, we clarify the distinction between direct and inverse first-order phase transitions in terms of the sign of the generalised pseudotrace, $\alpha_{\vartheta}$ (cf. Eq.˜7). In Ref. [76] it was observed that the quantity that most clearly distinguishes the different hydrodynamic solutions is a suitable generalisation of the trace anomaly of the stress-energy tensor, namely generalized pseudotrace $\alpha_{\vartheta}$. To make contact with more familiar thermodynamic quantities, it was shown that this object coincides with the standard thermodynamic latent heat $L\equiv\Delta w(T_{c})$ when evaluated at the critical temperature. It is therefore natural to interpret $\alpha_{\vartheta}$ as the effective latent heat of the phase transition computed at the nucleation temperature rather than at the critical one. Then:
  • Direct PT ($\alpha_{\vartheta}>0$). The transition releases energy into the environment. The injected energy drives an outward pressure gradient in the plasma, sourcing fluid outflows from the advancing bubble wall, going “away” as seen from the center of the bubble.
  • Inverse PT ($\alpha_{\vartheta}<0$). The transition absorbs energy from the environment. Energy is drawn from the surrounding plasma into the bubble interface, which tends to establish inward pressure gradients and correspondingly favors fluid motion toward the center of the bubble.
In this description, “direct” versus “inverse” refers solely to the sign of $\alpha_{\vartheta}$: positive $\alpha_{\vartheta}$ implies energy release; negative $\alpha_{\vartheta}$ implies energy uptake. Because inverse transitions require extracting heat from the medium, they are intrinsically thermal processes and cannot occur at $T=0$.
It is worth to briefly comment on the conceptual difference between inverse and symmetry-restoring first-order phase transitions. In a symmetry-restoring transition, the system evolves from a broken-symmetry phase to a symmetric one; from the symmetry point of view, it does not matter whether this happens upon cooling or heating of the plasma, as the notion of “symmetry restoration” only tracks the order parameter. By contrast, an inverse phase transition is defined in hydrodynamical terms and is, in principle, independent of whether a symmetry is being restored or broken, and of whether the transition proceeds during cooling or heating. What matters is instead the sign of the generalised pseudotrace introduced above, which fixes the direction of the fluid flow: in the inverse case, the plasma is drawn into the bubble rather than being pushed outwards. Examples of symmetry-breaking inverse phase transitions occurring upon cooling have been presented in Ref. [76], while examples of inverse symmetry-restoring phase transitions taking place during heating have been discussed in Ref. [65, 66]. Notice that the phase transition presented in Ref. [77] was strictly speaking a symmetry-restoring direct phase transition.
This reversed fluid flow has profound implications for the dynamics of the transition. The hydrodynamic solutions that describe the plasma motion, such as detonations, deflagrations, and hybrids, are qualitatively modified, as the pressure and velocity profiles now correspond to an inward rather than outward propagation. From the point of view of the energy budget, ordinary (“direct”) first-order phase transitions are basically fueled by the $T=0$ vacuum energy difference between the two phases. For inverse transitions, instead, the situation is qualitatively different. At the nucleation temperature, the $T=0$ vacuum energy difference has the opposite sign, and needs to be interpreted as a product of the transition rather than part of the initial energy budget. In this sense, it is the thermal contribution from the plasma that drives the bubble expansion.
@@ II.1 Hydrodynamics of inverse phase transitions
beyond which either the detonation or the deflagration branch is selected and the composite construction ceases to satisfy the matching and causality constraints. By definition, [^3] First introduced in Ref. [75], $v_{J,\rm inv}$ marks the transition between inverse detonation and inverse hybrid regimes and can be viewed as the wall speed of the fastest inverse detonation (or, equivalently, the slowest inverse hybrid). $v_{J,\rm inv}$ is the boundary value at which the rarefaction and compression attachments become sonic and the hybrid solution bifurcates into a single-wave configuration.
@@ III The velocity in the LTE regime
In principle, the bubble wall velocity is not fixed by hydrodynamical considerations alone, and requires input from the microphysics of the plasma. This can be seen by counting the number of quantities that describe the transition, four, $T_{+},T_{-},v_{+},v_{-}$, compared to the two conservation equations of the stress-energy tensor, Eq.˜3. Hence, even for a given nucleation temperature (which is determined by the nucleation rate), one degree of freedom remains undetermined, which we typically take to be the wall velocity $\xi_{w}$. Determining $\xi_{w}$ from first principles requires a detailed study of the microscopic friction exerted by the plasma on the expanding wall.
[…]
Nevertheless, for exploratory purposes, we will adopt the standard LTE treatment, which allows one to capture the qualitative behavior of the wall and obtain a reasonable order-of-magnitude estimate of its velocity. This is particularly useful for mapping the parameter space of novel scenarios, such as inverse phase transitions [75, 76], where a complete out-of-equilibrium computation is not yet available.
@@ III.1 LTE framework
With this equation at hand, we can study the pressure for the direct and the inverse phase transitions. First of all, we can study this relation in the case of a just-nucleated bubble. The bubble is taken to be nucleated at rest with a vanishing wall velocity, $\xi_{w}\to 0$, and correspondingly $v_{+}=v_{-}\to 0$, consequently we have that
@@ III.2 LTE pressure for direct phase transitions
Figure 4: Phase diagram of the LTE wall velocity in the $(\alpha_{N},\Psi)$ plane. Coloured squares show the numerical wall speed. The solid black boundary is the full $\mathcal{P}_{\mathrm{LTE}}=0$ contour: its straight left segment is the static-wall limit $\mathcal{P}_{\mathrm{LTE}}(\xi_{w}=0)=0$ (a newly nucleated bubble does not expand). In the direct region, the curved branch marks the onset of runaway, equivalently $\mathcal{P}_{\mathrm{LTE}}(\xi_{w}=\xi_{J})<0$; in the inverse region, it is set by the slowest hybrid just past the inverse Jouguet speed, $\mathcal{P}_{\mathrm{LTE}}(\xi_{w}=\xi_{J,\mathrm{inv}}^{+})<0$. The green band highlights parameters satisfying the runaway criterion. The dotted vertical line at $\alpha_{N}\simeq-0.082$ indicates the point at which the kinematically forbidden region opens up, and the slowest inverse hybrid solution is always a stable solution.
@@ III.3 The LTE pressure for inverse phase transitions
As such, this boundary is expressed in terms of $\alpha_{+}$. To express it in terms of $\alpha_{N}$, we then rescale $\alpha_{+}\to\alpha_{N}\,\frac{\alpha_{+,\rm slowest}}{\alpha_{N,\rm slowest}}\,$, and substitute into Eq.˜30 to obtain the black curve delineating the transition between LTE‐velocity solutions and true runaway.
@@ IV Sound-shell model in an expanding Universe
In this section, we compute the gravitational-wave signal from first-order phase transitions with the explicit goal of disentangling direct from inverse dynamics. Our strategy is to model the acoustic stage within the sound-shell framework, in which the velocity field produced by the expanding bubbles is described as a superposition of spherically symmetric shells whose overlap sources tensor perturbations on a Friedmann background. Over the past years the framework has been thoroughly developed and applied to phase transitions of weak to intermediate strength. Starting from its original formulation and first quantitative predictions [85, 86], it has been used extensively in observational studies [91, 92, 93, 94, 95], and subsequently generalised to account for the role of the cosmological expansion in shaping the build-up and time coherence of the source [87] as well as for non-standard equations of state [89]. Here we adopt the formulation that clarifies the infrared momentum dependence and the emergence of intermediate spectral slopes [88], which we implement to track how the hydrodynamic differences between direct and inverse transitions propagate into the shape and normalization of the spectrum.
@@ V.1 Spectral angles: from pure shape to amplitude-aware comparisons
We highlight a distinctive low-velocity effect: the modified spectral angle tends to flatten as the wall speed decreases. Tracing this back to the energy budget, we find that the multi-bubble kinetic fraction $\Omega_{K}$ also flattens in the $\xi_{w}\!\to\!0$ limit. A steeper suppression might have been expected by considering the single-bubble efficiency only, which is known to drop rapidly at small $\xi_{w}$ [75]. As we argue in Appendix˜B, this is due to the fact that the overlap of many bubbles at collision compensates for the poor conversion of the isolated expanding bubbles.
@@ Acknowledgements.
It is a pleasure to thank Mark Hindmarsh, Simona Procacci and Alberto Roper Pol for helpful discussions, as well as Wenyuan Ai and Benoit Laurent for fruitful comments on the manuscript. EM and GB acknowledge support from the Spanish Research Agency (Agencia Estatal de Investigación, MCIN/AEI/10.13039/501100011033) via the IFT Severo Ochoa Center of Excellence grant CEX2020-001007-S. In addition, EM is supported through the grants CNS2023-14453 and PID2022-137127NB-I00 funded by MCIN/AEI/10.13039/501100011033, the European Union NextGenerationEU/PRTR and ESF+, and GB is supported by the grant CNS2023-145069 funded by MICIU/AEI/10.13039/501100011033 and by the European Union NextGenerationEU/PRTR. M.V. is funded by the European Union (ERC, HoloGW, Grant Agreement No. 101141909). Views and opinions expressed are, however, those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. M.V. also acknowledges financial support from Grant CEX2024-001451-M funded by MICIU/AEI/10.13039/501100011033, from Grant No. PID2022-136224NB-C22 from the Spanish Ministry of Science, Innovation and Universities, and from Grant No. 2021-SGR-872 funded by the Catalan Government and by the “Excellence of Science - EOS” - be.h project n.30820817, and by the Strategic Research Program High-Energy Physics of the Vrije Universiteit Brussel. SB is supported by the Deutsche Forschungsgemeinschaft under Germany’s Excellence Strategy - EXC 2121 Quantum universe - 390833306.